OpenStudy (jacksonjrb):

Write the equation for a circle with the following characteristics: -the center is on the line whose equation is 3y+12=2x -is tangent to both the x and y axis Only an algebraic solution will be accepted

OpenStudy (jacksonjrb):

y=2/3x-4 is the standard form of the equation for the line

OpenStudy (welshfella):

so the circle passes though the points (0,y1) and (x1,0) and its centre is at ( (2/3x2 - 4, y2)

OpenStudy (jacksonjrb):

i graphed the line and it crosses through points (0,6) and (-4,0)

OpenStudy (jacksonjrb):

that means the center is (-3,2) i believe. i dont know how to do that algebraically though

OpenStudy (welshfella):

right and the radius is also equal to y2

OpenStudy (jacksonjrb):

how could i show this algebraically

OpenStudy (welshfella):

( x - (2/3 x2 - 4)^2 + (y - r)^2 = r^2

OpenStudy (welshfella):

this is an interesting one but i haven't got my head around it yet. Unfortunately i have to go right now. sorry

OpenStudy (jacksonjrb):

ok thanks

OpenStudy (welshfella):